# Dynamic Integration of Task-Specific Adapters for Class Incremental Learning

### v1

@@ Abstract
Non-exemplar class Incremental Learning (NECIL) enables models to continuously acquire new classes without retraining from scratch and storing old task exemplars, addressing privacy and storage issues. However, the absence of data from earlier tasks exacerbates the challenge of catastrophic forgetting in NECIL. In this paper, we propose a novel framework called Dynamic Integration of task-specific Adapters (DIA), which comprises two key components: Task-Specific Adapter Integration (TSAI) and Patch-Level Model Alignment. TSAI boosts compositionality through a patch-level adapter integration strategy, which provides a more flexible compositional solution while maintaining low computation costs. Patch-Level Model Alignment maintains feature consistency and accurate decision boundaries via two specialized mechanisms: Patch-Level Distillation Loss (PDL) and Patch-Level Feature Reconstruction method (PFR). Specifically, the PDL preserves feature-level consistency between successive models by implementing a distillation loss based on the contributions of patch tokens to new class learning. The PFR facilitates accurate classifier alignment by reconstructing old class features from previous tasks that adapt to new task knowledge. Extensive experiments validate the effectiveness of our DIA, revealing significant improvements on benchmark datasets in the NECIL setting, maintaining an optimal balance between computational complexity and accuracy. The full code implementation will be made publicly available upon the publication of this paper.
Introduction
@@
Figure 1: (a). Shared parameter space leads to task interference. (b). High computation costs characterize current PET-based methods. (c). Gaussian-based feature reconstruction may have a significant deviation from the actual feature distribution. $c_{i}$ indicates the actual feature distribution of old classes, $\hat{c_{i}}$ indicates the feature distribution generated by Gaussian Sampling.
@@
Class Incremental Learning (CIL) has gained increasing attention within the expansive field of artificial intelligence research (Li et al. 2024b; Zhu et al. 2021b; Wang et al. 2022c; Kurniawan et al. 2024; Li et al. 2024a; Wang et al. 2023a; Bonato et al. 2024; Zhai et al. 2024a), offering the potential for models to continuously learn new knowledge while maintaining the knowledge of previously encountered old classes. Replay-based CIL approaches (Zhu et al. 2021b; Li and Hoiem 2018a; Wang et al. 2022a; Zhou et al. 2023; Rebuffi et al. 2017; Yan, Xie, and He 2021) require the storage of exemplars from previous tasks, which presents challenges in terms of memory limitations and privacy concerns. Benefiting from the advances in Pre-Trained Models (PTMs) (Dosovitskiy et al. 2020), Non-Exemplar Class Incremental Learning (NECIL) has become a promising alternative, enabling the incremental acquisition of knowledge without the need to maintain a buffer of old class exemplars.
@@
Despite the inherent ability of PTMs to produce generalizable features, which has led to superior performance in the NECIL setting, NECIL remains challenging due to two primary factors. First, As shown in Fig.1 (a), some studies (Gao et al. 2023; Zhang et al. 2023; Zhou et al. 2024a) that tune a shared parameter space for all tasks unable to segregate parameters for each task. The mixing of parameters leads to task interference, which in turn causes catastrophic forgetting. Particular PET-based methods (Wang et al. 2022b, c; Smith et al. 2023; Gao, Cen, and Chang 2024; Kurniawan et al. 2024) isolate task parameters by designing task-specific prompts. However, these methods establish a strong correlation between prompt selection and the class tokens. This tight coupling impedes the model from leveraging intact old task parameters to reconstruct knowledge without old exemplars, undermining the model’s capacity for effective knowledge retention and reproduction in NECIL scenarios. Moreover, these methods, requiring multiple forward propagations, significantly increase computational costs, as shown in Fig.1 (b). We summarize the above issues as the compositionality deficiency.
Second, the absence of exemplars from old tasks prevents the model from applying replay techniques to preserve old task knowledge. Currently, PET-based methods (Wang et al. 2022b, c; Smith et al. 2023; Gao, Cen, and Chang 2024) neglect the importance of maintaining consistency between incrementally trained models. This neglect causes the model to easily overfit new tasks, focusing solely on information relevant to new tasks while discarding others that may be critical for old tasks. It also fails to adapt classifiers to the decision boundary changes introduced by new tasks. There are prototype-based techniques (Zhang et al. 2023; Zhu et al. 2021b; Wang et al. 2023b) that employ the Gaussian distribution to model the distribution of old task features and align the classifier through the generated features. However, as shown in Fig.1 (c), the features created by Gaussian sampling increasingly deviate from the actual features of real images as new tasks are learned. We attribute the above issues to the model alignment deficiency.
@@
To mitigate the above two challenges, we introduce the Dynamic Integration of task-specific Adapters (DIA) framework, which comprises two main components: Task-Specific Adapter Integration (TSAI) and Patch-Level Model Alignment. Specifically, the TSAI module is designed to boost compositionality through a patch-level adapter integration strategy, which provides a more flexible compositional solution while maintaining low computation costs. Furthermore, we demonstrate the knowledge retention and reconstruction potential of TSAI on a parameter factorization basis. Building these TSAI’s capabilities, we propose a patch-level model alignment strategy to maintain feature consistency and accurate decision boundaries. The proposed model alignment strategy integrates two fundamental parts: 1) The Patch-Level Distillation Loss (PDL) maintains feature-level consistency between the old and new models by introducing a distillation loss based on the normalized distance of the patch tokens. PDL evaluates the contribution of patch tokens to the new task learning and penalizes feature drift in non-contributory tokens to maintain feature consistency. 2) The Patch-Level Feature Reconstruction (PFR) employs a normalized distance difference to retrieve patch tokens associated with previously acquired task knowledge. These tokens, combined with old class prototypes, are then used to reconstruct old class features that adapt to the newly learned tasks.
@@
Comprehensive experiments on four benchmark datasets demonstrate the superiority of the proposed DIA method, which achieves state-of-the-art (SOTA) performance in the NECIL setting. Moreover, as shown in Fig.1 (a), our DIA maintains up to a 90% decrease in computational complexity while maintaining SOTA performance.
The contributions can be summarized as follows:
@@
  • We present a patch-level model alignment strategy based on the knowledge retention and reconstruction capability of TSAI, incorporating PDL to maintain feature consistency and PFR to reconstruct old class features that adapt to new knowledge, improving classifier alignment.
@@
  • We conduct extensive experiments across four benchmarks, where the proposed DIA achieves state-of-the-art performance while maintaining an optimal balance between computational complexity and accuracy.
Related Work
@@
Parameter-Efficient Tuning: As an efficient alternative to full fine-tuning, Adapter tuning (Houlsby et al. 2019) was initially introduced to efficiently transfer large pre-trained models to downstream tasks in NLP tasks. Afterward, methods such as Prompt-Tuning (Lester, Al-Rfou, and Constant 2021) and Prefix-Tuning (Li and Liang 2021) adapt models by inserting learnable tokens explicitly tailored to the new tasks. Following the success of Vision Transformers (Dosovitskiy et al. 2020; Liu et al. 2021), PET methods have been adapted for visual transfer learning. Notable examples include Visual Prompt Tuning (VPT) (Jia et al. 2022) and AdapterFormer (Chen et al. 2022), which apply PET techniques to vision tasks. These methods achieve comparable or superior performance to full fine-tuning while maintaining efficiency. In this paper, we propose a Dynamic Integration of task-specific Adapters framework for the NECIL problem based on PET methods (Chen et al. 2022; Houlsby et al. 2019).
Non-exemplar Class Incremental Learning NECIL approaches address privacy and memory concerns by eliminating the need for old task exemplars and employing various techniques such as regularization (Li and Hoiem 2018b; Huang et al. 2024; Yu et al. 2020), augmentation (Zhu et al. 2021a, b; Kim, Park, and Han 2024), and model rectification-based (Wang et al. 2023b; Zhu et al. 2021b) methods to maintain model performance across tasks. Currently, PTM-based methods that sequentially adjust the PTM to stream data with new classes are a promising direction for NECIL. Most methods (Wang et al. 2022c, b; Smith et al. 2023; Gao, Cen, and Chang 2024; Roy et al. 2024) focus on the instance-level prompt selection to segregate task parameters. LAE (Gao et al. 2023) and ADAM (Zhou et al. 2024a) propose unified frameworks for PET methods by model ensemble. SLCA (Zhang et al. 2023) extends the Gaussian modeling (Zhu et al. 2021b) of old task features to rectify classifiers. EASE (Zhou et al. 2024b) utilizes adapters to extract task-specific features through multiple forward propagations.
@@
Figure 2: Illustration of DIA. (a) Task-Specific Adapter Integration. For incremental task $t$, we learn a task adapter $\mathcal{A}^{t,b}$ and a task signature vector $\bm{\tau}^{t,b}\in\mathcal{R}^{d}$ at each transformer block $b$. Each token is routed to the relevant task adapters through the signature vectors, processed independently, and combined into an integrated, task-informed output. (b) Patch-level Distillation. We encourage feature drift in the patch tokens that contribute to new task learning while penalizing others that do not. (c) Patch-level Feature Reconstruction. We identify patch tokens in the patch tokens that are related to old class knowledge and integrate them with the old class prototype $\bm{\mu}_{k}^{t-1}$ to reconstruct old class feature $\hat{\bm{\mu}}_{k}^{t-1}$.
Problem Setup
@@
In this paper, we focus on the NECIL setting, which prohibits storing exemplars from previous tasks. NECIL involves sequentially learning a set of $T$ tasks, denoted as $\{\mathcal{T}^{t}\}_{t=1}^{T}$, where each task $\mathcal{T}^{t}=\{D^{t},C^{t}\}$ consists of a current training set $D^{t}=\{(I^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$ and a class label set $C^{t}=\{c^{t}_{k}\}_{k=1}^{M^{t}}$. In this context, $I^{t}_{i}$ is the input image, $y^{t}_{i}$ is the class label for the $i$-th image belonging to $C^{t}$, $N^{t}$ represents the number of images, and $M^{t}$ denotes the number of classes in $C^{t}$. There is no overlap between the classes of different tasks, meaning $\forall i,j,\ C^{i}\cap C^{j}=\varnothing$. After training on a task $\mathcal{T}^{t}$, the model is evaluated on all the classes encountered so far, $C^{1:t}=C^{1}\cup C^{2}\cup\cdots\cup C^{t}$.
Methodology
Overview
@@
As illustrated in Fig.2, we propose the Dynamic Integration of task-specific Adapters (DIA) framework for the NECIL scenario. At incremental task $t$, we introduce a task-specific adapter $\mathcal{A}^{t,b}$ parallel to the MLP layer and a task signature vector $\bm{\tau}^{t,b}$ into each transformer block $b$. As shown in Fig.2 (a), each image token is routed by the signature vectors $[\bm{\tau}^{i,b}]_{i=1}^{t}$ to relevant adapters. These task-specific adapters independently process the input token, and their outputs are merged using scalars determined by the task signature vectors, yielding an integrated output. At task $t$, only $\mathcal{A}^{t,b}$ and $\bm{\tau}^{t,b}$ are trainable. Leveraging TSAI’s capacity for knowledge retention and reconstruction, along with the rich information embedded in patch tokens, we introduce a patch-level model alignment mechanism at both the feature and classifier levels. Firstly, as shown in Fig.2 (b), we compute the Patch-Level Distillation Loss (PDL) based on the feature drift between the patch tokens $p^{n},p^{o}$ obtained from the model trained on task $t$ and task $t-1$. PDL penalizes the feature drift in patch tokens that do not contribute to the new task learning to preserve old task knowledge. Secondly, we propose a Patch-level Feature Reconstruction (PFR) method to reconstruct old class features without exemplars, as shown in Fig.2 (c). We identify patch tokens related to old task knowledge and integrate them with the old task prototypes. The reconstructed features are utilized to calibrate the decision boundaries of the classifier, adapting to the changes in feature distribution.
Task-Specific Adapters Integration
@@
The proposed TSAI aims to provide a more flexible compositional solution by employing a patch-level adapter integration strategy. To achieve this goal, we learn a task-specific adapter $\mathcal{A}^{t,b}$ and a task signature vector $\bm{\tau}^{t,b}\in R^{d}$ for each incremental task, where $d$ indicates feature dimension, $t$ represents the incremental task id, and $b$ represents the block index. For convenience, we omit block index $b$, as the computation is identical across transformer blocks.
@@
Specifically, the task-specific adapter $\mathcal{A}^{t}$ is a bottleneck module comprising a down-projection layer $W_{\text{down}}\in\mathbb{R}^{d\times r}$, an up-projection layer $W_{\text{up}}\in\mathbb{R}^{r\times d}$. This bottleneck module extends and adjusts the feature space by modifying the MLP output through the residual connection via a scaling vector $\mathbf{s}^{t}$ obtained by $\bm{\tau}^{t}$. For an input $\mathbf{X}={\left[{\mathbf{p}_{j}}^{\top}\right]}_{j=0}^{L}\in R^{(L+1)\times d}$ with $L+1$ image tokens, where $\mathbf{p}_{0}\in\mathbb{R}^{d}$ indicates the class token and $\mathbf{p}_{j}\in\mathbb{R}^{d},j>1$ indicate the patch token. The task signature vector $\bm{\tau}^{t}$ assigns scalar $s_{j}^{t}$ to image token $\mathbf{p}_{j}$ to evaluate its task relevance:
$\mathbf{\tilde{p}}_{j}^{t}$ $=s_{j}^{t}\mathcal{A}^{t}(\mathbf{p}_{j})=s_{j}^{t}\text{ReLU}(\mathbf{p}_{j}W_{\text{down}})W_{\text{up}},$ (1)
$\quad\quad s_{j}^{t}$ $=<\bar{\mathbf{p}}_{j},\bar{\bm{\tau}}^{t}>,\mathbf{s}^{t}=\left[s_{j}^{t}\right]_{j=0}^{L},$ (2)
@@
where, $\bar{\mathbf{p}}_{j},\bar{\bm{\tau}}^{t}$ represent normalized tensors $\bar{\mathbf{p}}=\mathbf{p}/\|\mathbf{p}\|_{2}$, $\bar{\bm{\tau}}^{t}=\bm{\tau}^{t}/\|\bm{\tau}^{t}\|_{2}$, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > represent the dot production.
@@
Afterwards, the output feature $\mathbf{o}_{j}$ for token $\mathbf{p}_{j}$ using one adapter $\mathcal{A}^{t}$ can be calculated as follows:
$\quad\quad\mathbf{o}_{j}$ $=\text{MLP}(\mathbf{p}_{j})+\mathbf{\tilde{p}}_{j}^{t}+\mathbf{\mathbf{p}_{j}}$
@@
For incremental task $t>1$, the output feature $\mathbf{o}_{j}$ of TSAI is integrated from multiple task-specific adapters $\{\mathcal{A}^{i}\}_{i=1}^{t}$. We use the softmax operation to normalize the scaling vector $[s_{j}^{i}]_{i=1}^{t}$ for patch token $\mathbf{p}_{j}$, allowing the model to combine the contributions from different adapters while maintaining a balanced and stable output.
$\mathbf{o}_{j}=\text{MLP}(\mathbf{p}_{j})+\sum_{i=1}^{t}{\hat{s}_{j}^{i}\mathcal{A}^{i}(\mathbf{p}_{j})}+\mathbf{p}_{j},$ (4)
$\mathbf{O}={\left[\mathbf{o}_{j}^{\top}\right]}_{j=0}^{L},\quad[\hat{s}_{j}^{i}]_{i=1}^{t}=\text{softmax}([s_{j}^{i}]_{i=1}^{t}),$ (5)
where $\mathbf{O}$ is the output of TSAI for input $\mathbf{X}$.
Analysis of knowledge of retention and reproduction: We conducted an in-depth analysis of the parameter space for each task-specific adapter, using matrix factorization to demonstrate that our proposed TSAI module can effectively preserve and reproduce knowledge from previous tasks without needing exemplars from old tasks. In particular, we start with no activation function to describe the knowledge retention mechanism under linear conditions and then extend to nonlinear conditions. We provide a more detailed analysis in the supplementary material.
To avoid confusion, assume the input token is $\mathbf{p}\in\mathbb{R}^{m}$ and the adapter weight is $W=W_{\text{down}}W_{\text{up}}\in\mathbb{R}^{m\times n}$, with a matrix rank of $r$. The weight matrix can be decomposed through SVD without information loss:
@@
$\mathbf{U}^{\top}$ $=\left[{\mathbf{u}_{i}}^{\top}\right]_{i=1}^{r}\in\mathbb{R}^{r\times m},\mathbf{V}={\left[{\mathbf{v}_{i}}^{\top}\right]_{i=1}^{r}}\in\mathbb{R}^{r\times n},$ (7)
where $\text{diag}(\sigma)$ is a diagonal matrix with singular values $\sigma_{i}$ on the diagonal. $\mathbf{U}$ is an $m\times r$ orthogonal matrix, with the singular vectors $\mathbf{u}_{i}$ as its columns. $\mathbf{V}$ is an $r\times n$ orthogonal matrix, with the singular vectors $\mathbf{v}_{i}^{\top}$ as its rows.

### v2

@@ Abstract
Non-exemplar Class Incremental Learning (NECIL) enables models to continuously acquire new classes without retraining from scratch and storing old task exemplars, addressing privacy and storage issues. However, the absence of data from earlier tasks exacerbates the challenge of catastrophic forgetting in NECIL. In this paper, we propose a novel framework called Dynamic Integration of task-specific Adapters (DIA), which comprises two key components: Task-Specific Adapter Integration (TSAI) and Patch-Level Model Alignment. TSAI boosts compositionality through a patch-level adapter integration strategy, aggregating richer task-specific information while maintaining low computation costs. Patch-Level Model Alignment maintains feature consistency and accurate decision boundaries via two specialized mechanisms: Patch-Level Distillation Loss (PDL) and Patch-Level Feature Reconstruction (PFR). Specifically, on the one hand, the PDL preserves feature-level consistency between successive models by implementing a distillation loss based on the contributions of patch tokens to new class learning. On the other hand, the PFR promotes classifier alignment by reconstructing old class features from previous tasks that adapt to new task knowledge, thereby preserving well-calibrated decision boundaries. Comprehensive experiments validate the effectiveness of our DIA, revealing significant improvements on NECIL benchmark datasets while maintaining an optimal balance between computational complexity and accuracy.
Yihong Gong and Shaokun Wang are the corresponding authors.
1 Introduction
@@
Figure 1: (a) Shared parameter space leads to task interference. (b) High computation costs characterize current PET-based methods. (c) Gaussian-based feature reconstruction may have a significant deviation from the actual feature distribution. $c_{i}$ indicates the actual feature distribution of old classes, $\hat{c_{i}}$ indicates the feature distribution generated by Gaussian Sampling.
@@
Class Incremental Learning (CIL) has gained considerable attention within the expansive field of artificial intelligence research [29, 51, 39, 18, 20, 28, 15, 3, 36, 1, 44], offering the potential for models to continuously learn new knowledge while maintaining the knowledge of previously encountered old classes. Replay-based CIL approaches [51, 24, 35, 46, 30, 40] require the storage of exemplars from previous tasks, which presents challenges in terms of memory limitations and privacy concerns. Benefiting from the advances in Pre-Trained Models (PTMs) [4], Non-Exemplar Class Incremental Learning (NECIL) has emerged as a promising alternative, enabling the incremental acquisition of knowledge without the need to maintain a buffer of old class exemplars.
@@
Despite the inherent ability of PTMs to produce generalizable features, which has led to superior performance in the NECIL setting, NECIL remains challenging due to two primary factors. First, as shown in Fig.1 (a), studies [5, 45, 48] that tune a shared parameter space for all tasks are unable to segregate parameters for each task. The mixing of parameters leads to task interference, which in turn causes catastrophic forgetting. Particular PET-based methods [38, 39, 33, 7, 18] isolate task parameters by designing task-specific prompts. However, these methods establish a strong correlation between prompt selection and the class tokens. This tight coupling neglects the rich semantic information embedded in patch tokens and the task-specific knowledge preserved in the intact old task parameters, thereby undermining the model’s capacity for effective knowledge retention and reproduction in NECIL scenarios. Moreover, these methods, requiring multiple forward propagations, increase computational costs, as shown in Fig.1 (b). We summarize the above issues as the compositionality deficiency.
Second, the absence of exemplars from previous tasks prevents the model from utilizing replay techniques, thereby posing significant challenges in maintaining feature consistency and preserving accurate decision boundaries. Current PET-based methods [38, 39, 33, 7] overlook the necessity of maintaining consistency across incrementally trained models. In incremental learning scenarios, the representations of old task samples will change due to the incorporation of new task knowledge, and this neglect ultimately leads to feature drift in the old task samples. Additionally, these approaches are unable to adapt classifiers to the decision boundary changes introduced by new tasks. Regularization-based approaches [43, 37, 51, 52, 22] that impose constraints on class tokens to maintain feature consistency often result in overly restrictive regularization, hindering the learning for new tasks. Prototype-based approaches [45, 51, 37] leverage Gaussian distributions to generate previous task features and then align classifiers. However, it can lead to inaccurate decision boundaries, as features generated by Gaussian sampling progressively diverge from the actual features as new tasks are introduced, as shown in Fig.1 (c). We attribute the above issues to the model alignment deficiency.
@@
To mitigate the above two challenges, we introduce the Dynamic Integration of task-specific Adapters (DIA) framework, which comprises two components: Task-Specific Adapter Integration (TSAI) and Patch-Level Model Alignment. Specifically, the TSAI module is designed to boost compositionality through a patch-level adapter integration strategy, which provides a more flexible compositional solution while maintaining low computation costs. Moreover, we demonstrate the knowledge retention and reconstruction potential of TSAI on a parameter factorization basis. Building upon these capabilities, we propose a patch-level model alignment strategy to preserve feature consistency and maintain accurate decision boundaries across tasks. This strategy integrates two fundamental parts: 1) The Patch-Level Distillation Loss (PDL) ensures feature-level consistency between the old and new models by introducing a distillation loss that regulates the feature shift of patch tokens. PDL evaluates the contribution of patch tokens to the new task learning and then penalizes feature drift in non-contributory tokens to maintain feature consistency. 2) The Patch-Level Feature Reconstruction (PFR) retrieves patch tokens that encapsulate knowledge from prior tasks and combines them with prototypes to reconstruct old class features aligned with the newly learned tasks. These reconstructed features facilitate classifier alignment, ensuring the maintenance of an accurate decision boundary throughout the incremental process.
@@
Comprehensive experiments on four NECIL benchmark datasets demonstrate that our DIA method achieves state-of-the-art (SOTA) performance. Moreover, as shown in Fig.1 (b), our DIA maintains up to a 90% decrease in computational complexity while maintaining SOTA performance. In summary, the main contributions include:
@@
  • We present a patch-level model alignment strategy based on TSAI: PDL maintains feature-level consistency by regulating feature drift in non-contributory patch tokens. PFR improves classifier alignment by reconstructing old class features aligned with newly learned tasks.
@@
  • Experimental results across four benchmarks demonstrate that our DIA achieves SOTA performance while maintaining an optimal balance between computational complexity and accuracy.
2 Related Work
@@
Parameter-Efficient Tuning: As an efficient alternative to full fine-tuning, Adapter Tuning [11] was initially introduced to efficiently transfer large pre-trained models to downstream tasks in NLP tasks. Afterward, methods such as Prompt-Tuning [19, 21] adapt models by inserting learnable tokens explicitly tailored to the new tasks. Following the success of Vision Transformers [4, 26], PET methods have been adapted for visual transfer learning. Notable examples include Visual Prompt Tuning (VPT) [13] and AdapterFormer [2], which apply PET techniques to vision tasks. These methods achieve comparable or superior performance to full fine-tuning while maintaining efficiency. In this paper, we propose a patch-level task adapter integration strategy that enables patch tokens to aggregate richer task information with greater flexibility while maintaining low computational costs.
@@
Figure 2: Illustration of DIA. (a) Task-Specific Adapter Integration. For incremental task $t$, we learn a task adapter $\mathcal{A}^{t,b}$ and a task signature vector $\bm{\tau}^{t,b}\in\mathcal{R}^{d}$ at each transformer block $b$. Each token is routed to the relevant task adapters through the signature vectors, processed independently, and combined into an integrated, task-informed output. (b) Patch-level Distillation. We promote feature drift in patch tokens that contribute to new task learning while penalizing those that do not, thus regulating the feature shift associated with old tasks. (c) Patch-level Feature Reconstruction. We identify patch tokens that are related to old class knowledge and integrate them with the old class prototype $\bm{\mu}_{k}^{t-1}$ to reconstruct old class feature $\hat{\bm{\mu}}_{k}^{t-1}$ aligned with new tasks.
Non-exemplar Class Incremental Learning NECIL approaches address privacy and memory concerns by eliminating the need for old task exemplars and employing various techniques such as regularization [23, 12, 42], augmentation [50, 51, 14], and model-alignment strategies [27, 8, 41, 37, 51] to maintain model performance across tasks. Particularly, SDC [41] calculates class prototype shifts using feature gaps. LDC [8] uses a forward projector to learn mappings from old to new feature spaces. Currently, PTM-based methods that sequentially adjust the PTM to stream data with new classes are a promising approach for NECIL. Most methods [39, 38, 33, 7, 31] focus on the instance-level ([cls] token-based) prompt selection to segregate task parameters. LAE [5] and ADAM [48] propose unified frameworks for PET methods by model ensemble. C-ADA [6] leverages Continual Adapter Layers and a Scale and Shift Module to learn new tasks. SLCA [45] extends the Gaussian modeling [51] of old task features to rectify classifiers. EASE [49] utilizes adapters to extract task-specific features through multiple forward propagations. In this paper, the proposed DIA framework leverages TSAI to enhance compositionality and incorporates patch-level model alignment to mitigate feature shifts and decision boundary distortions.
3 Problem Setup
@@
NECIL involves sequentially learning a set of $T$ tasks, denoted as $\{\mathcal{T}^{t}\}_{t=1}^{T}$, where each task $\mathcal{T}^{t}=\{D^{t},C^{t}\}$ consists of a current training set $D^{t}=\{(I^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$ and a class label set $C^{t}=\{c^{t}_{k}\}_{k=1}^{M^{t}}$. In this context, $I^{t}_{i}$ is the input image, $y^{t}_{i}$ is the class label for the $i$-th image belonging to $C^{t}$, $N^{t}$ represents the number of images, and $M^{t}$ denotes the number of classes in $C^{t}$. There is no overlap between the classes of different tasks, meaning $\forall i,j,\ C^{i}\cap C^{j}=\varnothing$. After training on a task $\mathcal{T}^{t}$, the model is evaluated on all the classes encountered so far, $C^{1:t}=C^{1}\cup C^{2}\cup\cdots\cup C^{t}$. Notably, NECIL prohibits storing exemplars from old tasks.
4 Methodology
4.1 Overview
@@
Fig.2 shows an overview of our DIA framework. At incremental task $t$, we introduce a task-specific adapter $\mathcal{A}^{t,b}$ in parallel with the MLP layer, along with a task signature vector $\bm{\tau}^{t,b}$ into each transformer block $b$. As shown in Fig.2 (a), each image token is routed by the signature vectors $[\bm{\tau}^{i,b}]_{i=1}^{t}$ to relevant adapters. These task-specific adapters independently process the input token, and their outputs are merged using scalars determined by the task signature vectors, yielding an integrated output. At the task $t$, only $\mathcal{A}^{t,b}$ and $\bm{\tau}^{t,b}$ are trainable. Leveraging TSAI’s capacity for knowledge retention and reconstruction, along with the rich information embedded in patch tokens, we introduce a patch-level model alignment mechanism at both the feature and classifier levels. First, as shown in Fig.2 (b), we compute the Patch-Level Distillation Loss (PDL) based on the feature drift between the patch tokens $\mathbf{p}^{t}$ and $\mathbf{p}^{t-1}$, which are obtained from the model trained on task $t$ and task $t-1$, respectively. PDL penalizes the feature drift in patch tokens that do not contribute to the new task learning, thereby preserving the old task knowledge. Second, we propose a Patch-level Feature Reconstruction (PFR) method to reconstruct old class features without exemplars, as shown in Fig.2 (c). We identify patch tokens related to old task knowledge and integrate them with the old task prototypes. The reconstructed features are then used to recalibrate the decision boundaries of the classifier, adapting to the changes in feature distribution.
4.2 Task-Specific Adapters Integration
@@
Our TSAI aims to provide a more flexible compositional solution by employing a patch-level adapter integration strategy. To achieve this goal, we learn a task-specific adapter $\mathcal{A}^{t,b}$ and a task signature vector $\bm{\tau}^{t,b}\in R^{d}$ for each incremental task, where $d$ indicates feature dimension, $t$ represents the incremental task id, and $b$ represents the block index. For convenience, we omit block index $b$, as the computation is identical across transformer blocks.
@@
Specifically, the task-specific adapter $\mathcal{A}^{t}$ is a bottleneck module comprising a down-projection layer $\mathbf{W}_{\text{down}}\in\mathbb{R}^{d\times r}$, an up-projection layer $\mathbf{W}_{\text{up}}\in\mathbb{R}^{r\times d}$. This bottleneck module extends and adjusts the feature space by modifying the MLP output through the residual connection, utilizing a scaling vector $\mathbf{s}^{t}$ derived from $\bm{\tau}^{t}$. For an input $\mathbf{X}={\left[{\mathbf{p}_{j}}^{\top}\right]}_{j=0}^{L}\in R^{(L+1)\times d}$, consisting of $L+1$ image tokens, where $\mathbf{p}_{0}\in\mathbb{R}^{d}$ indicates the class token and $\mathbf{p}_{j}\in\mathbb{R}^{d},j>0$ denote the patch tokens. The task signature vector $\bm{\tau}^{t}$ assigns scalar $s_{j}^{t}$ to each image token $\mathbf{p}_{j}$ to evaluate its task relevance:
$\mathbf{\tilde{p}}_{j}^{t}=s_{j}^{t}\mathcal{A}^{t}(\mathbf{p}_{j})=s_{j}^{t}\text{ReLU}(\mathbf{p}_{j}\mathbf{W}_{\text{down}})\mathbf{W}_{\text{up}},$ (1)
$s_{j}^{t}=<\bar{\mathbf{p}}_{j},\bar{\bm{\tau}}^{t}>,\mathbf{s}^{t}=\left[s_{j}^{t}\right]_{j=0}^{L},$ (2)
@@
where $\bar{\mathbf{p}}_{j},\bar{\bm{\tau}}^{t}$ represent normalized tensors $\bar{\mathbf{p}}=\mathbf{p}/\|\mathbf{p}\|_{2}$, $\bar{\bm{\tau}}^{t}=\bm{\tau}^{t}/\|\bm{\tau}^{t}\|_{2}$, respectively, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > represents the dot product.
@@
Subsequently, the output feature $\mathbf{\hat{p}}_{j}$ for the token $\mathbf{p}_{j}$ using the adapter $\mathcal{A}^{t}$ can be calculated as follows:
$\mathbf{\hat{p}}_{j}$ $=\text{MLP}(\mathbf{p}_{j})+\mathbf{\tilde{p}}_{j}^{t}+\mathbf{\mathbf{p}_{j}}$
@@
For incremental task $t>1$, the output feature $\mathbf{\hat{p}}_{j}$ of TSAI is integrated from multiple task-specific adapters $\{\mathcal{A}^{i}\}_{i=1}^{t}$. We apply the softmax operation to normalize the scaling vector $[s_{j}^{i}]_{i=1}^{t}$ for the patch token $\mathbf{p}_{j}$, enabling the model to combine the contributions from different adapters while maintaining a balanced and stable output.
$\mathbf{\tilde{p}}_{j}^{t}=\sum_{i=1}^{t}{\hat{s}_{j}^{i}\mathcal{A}^{i}(\mathbf{p}_{j})},$ (4)
$\mathbf{\hat{p}}_{j}=\text{MLP}(\mathbf{p}_{j})+\mathbf{\tilde{p}}_{j}^{t}+\mathbf{p}_{j},$ (5)
$\mathbf{O}^{t}={\left[\mathbf{\hat{p}}_{j}^{\top}\right]}_{j=0}^{L},\quad[\hat{s}_{j}^{i}]_{i=1}^{t}=\text{softmax}([s_{j}^{i}]_{i=1}^{t}),$ (6)
where $\mathbf{O}^{t}$ is the output of TSAI for input $\mathbf{X}$.
Knowledge Retention and Reconstruction Analysis: We conduct an in-depth analysis of the parameter space for each task-specific adapter, using matrix factorization to demonstrate that our TSAI module can effectively preserve and reproduce knowledge from previous tasks without needing exemplars. In particular, we start with no activation function to describe the knowledge retention mechanism under linear conditions and then extend to nonlinear conditions. We provide a detailed analysis in the supplementary.
To avoid confusion, we define the input token as $\mathbf{p}\in\mathbb{R}^{d}$ and the adapter weight as $\mathbf{W}=\mathbf{W}_{\text{down}}\mathbf{W}_{\text{up}}\in\mathbb{R}^{d\times d^{{}^{\prime}}}$, with a matrix rank of $r$. The weight matrix can be decomposed through SVD [16] without information loss:
@@
$\mathbf{U}^{\top}$ $=\left[{\mathbf{u}_{i}}^{\top}\right]_{i=1}^{r}\in\mathbb{R}^{r\times d},\mathbf{V}={\left[{\mathbf{v}_{i}}^{\top}\right]_{i=1}^{r}}\in\mathbb{R}^{r\times d^{{}^{\prime}}},$ (8)
where $\text{diag}(\sigma)$ is a diagonal matrix with singular values $\sigma_{i}$ on the diagonal. $\mathbf{U}$ is an $d\times r$ orthogonal matrix, with the singular vectors $\mathbf{u}_{i}$ as its columns. $\mathbf{V}$ is an $r\times d^{{}^{\prime}}$ orthogonal matrix, with the singular vectors $\mathbf{v}_{i}^{\top}$ as its rows.